OpenStudy (chriscollins0097):

Suppose a normal distribution has a mean of 98 and a standard deviation of 6. What is P(x<_ 92)? A. 0.975 B. 0.84 C. 0.025 D. 0.16

OpenStudy (holsteremission):

\[\mathbb P(X\le92)=\mathbb P\left(\frac{X-98}{6}\le\frac{92-98}{6}\right)=\mathbb P(Z\le -1)\]where follows the standard normal distribution. You can use the empirical rule or a left-tailed \(z\) chart to look up the probability.

OpenStudy (loviy):

The work above is correct! To make it clearer, note that the formula used is: \[z=\frac{x-\mu}{\sigma} \] where \[\mu\] is the mean and is the standard deviation